### v1

# Can the 3 neutrino masses really be found using SN 1987A data?

Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remrkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, a theoretical linkage to possible CPT violation, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumotions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any masses $1eV/c^{2}$ would be inconsistent with upper limits from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A is too great for the method to work.
Keywords: SN 1987A, Neutrino, supernova, CPT violation, tachyon, KATRIN
I Introduction
About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A). Of those $10^{58}$ neutrinos only a couple of dozen were captured by Earth’s detectors. Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in our galaxy since neutrino detectors first existed. The early articles on the neutrino mass obtained from SN 1987A data mostly, with one exception,(Cowsik1988, ) set upper limits on its value. (Arnett1987, ; Burrows1988, ) The upper limits restriction is based on the almost universal belief that the differences in the three neutrino mass states are simply too small to use the data to find the individual masses themselves.
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) an extensive discussion of CPT violation, (2) a much improved statistical analysis, (3) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A.
I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
Many theorists have regarded quantum fields giving rise to physical tachyons as being either impossible or unlikely in light of the many theoretical difficulties they create. However, unobservable $m^{2}<0$ particles seem to be OK with theorists. Thus, for most theorists the idea that symmetry-breaking in the interaction of Higgs fields with massless gauge fields leads to the production of unobservable tachyons is not objectionable. The difficulties with “physical” tachyons having an $m^{2}<0$ that can be observed include their energy spectrum being unbounded from below, their frame-dependent and unstable vacuum state, and their noncovariant commutation rules. Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al. (Paczos2024, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues. Charles Schwartz is another theorist who has challenged all the reasons why tachyons are considered unphysical. (Schwartz2022, ).
I.2 The meaning of negative searches for tachyons
The question of whether $m^{2}<0$ and $v>c$ observable particles really exist is not theoretical, but strictly empirical. All reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect. (Ehrlich2022, ) Most physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category. A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable. A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA. However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is very close to zero, so we cannot tell if $m^{2}>0$ or $m^{2}<0$. Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal. Negative results for tachyonic neutrinos are not unlike negative results for searches for extraterestrials (SETI). In both cases, an appropriate response to a negative search would be “keep looking using more sensitive techniques.” Surprisinfly, however, if you ask an AI whether negative searches for tachyons make it more likely they don’t exist and then ask the same question about ET’s existence, you are illogically told yes for the former question and no for the latter.
I.3 The KATRIN experiment
The KATRIN experiment seeks to find the “effective mass” of the electron antineutrino $m^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}.$ Its best value as of 2025 is negative: $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the size of the uncertainty. Is it just a coincidence that the KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass value the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}?$ Ehrlich2015 Maybe it is a just a weird coincidence, but fortunately, as KATRIN takes more data the statistical uncertainty on the $m^{2}$ value now $0.108eV^{2}$ will be reduced. Whatever the final KATRIN results, recall that the experiment measures the effective mass of the electron neutrino and not the three individual masses comprising it, which is the primary interest of this paper.
II 3 Neutrino masses and SN 1987A
It is widely believed that we can set an upper limit on the neutrino masses based on oscillation experiments. These experiments yield values for separations in the mass squared of pairs of mass states: $\Delta m^{2}_{j,k}=m_{j}^{2}-m_{k}^{2},$ the two well-established ones being the atmospheric and solar $m^{2}$ splittings $\Delta m^{2}_{atm}$ and $\Delta m^{2}_{sol}.$
II.1 A true upper limit on Neutrino masses?
The three masses $m^{2}_{k}$ values are normally assumed to be no greater than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}\approx 0.0025eV^{2}.$ It is important to recognize this upper limit is based on a questionable assumption, namely the non-existence of any third observed oscillation frequency. In other words, since the two oscillation frequencies found for (1) $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ and (2) $\Delta m^{2}_{atm}$ are so close, they are empirically indistinguishable. If a third observed oscillation frequency did exist (perhaps the result of a fourth sterile neutrino) then the assumption of the upper limit $m_{k}^{2}<0.0025eV^{2}$ would be false. In that case, it is conceivable the neutrino masses are large enough to be measurable based on the neutrino travel time over a very large distance. Moreover, the best place to measure the three masses based on travel time might be a galactic supernova, provided the neutrino emissions from it are nearly simultaneous. No other source of a neutrino burst, even that produced by an atomic bomb, could come close to the precision made possible by the enormous travel distance to a supernova.
II.2 Supernova SN 1987A neutrinos
Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987. Those detectors (Kamiokande II, IMB, and Baksan) recorded bursts of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos. The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds. This lack of synch is partially removed in most analyses by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay
$\bar{\nu_{e}}+p\rightarrow n+e^{+}$ (1)
so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy. Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened. This late light arrival, of course does not imply that neutrinos are superlumnial. Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter. Thus, photons but not neutrinos are delayed by interactions with the supenova ejecta.
II.3 The 5 hour early LSD neutrino burst
In addition to the three detectors, there was a fourth small detector known as LSD ln 1987 which was located under Mont Blanc. LSD reported a burst of five neutrinos which occurred 484 min before the other three detectors’ bursts. This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival. If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones ones as well. (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists.
The five neutrinos observed during a 7 second time interval in LSD would if spurious represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months. In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors. However, the LSD neutrinos were all very low energy, and the other detectors could not have seen an LSD-like burst at that early time because of their higher energy thresholds. Thus, for example, the largest detector, Kamiokande II, had an energy threshold of 7 MeV, well above the energies of four of the five LSD neutrinos constituting the burst Another suspicious fact is that the main t=0 burst at 7:35 UT was not seen in LSD. However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst). Two final major problem with the LSD burst being associated with SN 1987A are (1) why the very low energies of all the 5 neutrinos comprising it ?, and (2 )why the virtual monochromaticity of the burst? In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five above objections to it are answerable, and we address the monochromaticity problem later.
III Finding $m^{2}$ of SN 1987A neutrinos
The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close to simultaneously. Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of $\emph{i}ndividual$ neutrinos. In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
If a photon’s speed en route to Earth is $c=D/T$ then a neutrino’s speed would be $v=D/(t+T)\approx c(1-t/T).$ Here t is the time delay of a neutrino relative to a photon. But since there can also be such a delay if a neutrino is emitted later than other neutrinos, so the definition of t only makes sense if almost all neutrinos were emitted during a very small time window. For the neutrino speed we have v/c=1t/T.v/c=1-t/T. and also $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}.$ Setting the two versions of $v/c$ equal
$v/c=1-t/T=1-m^{2}c^{4}/2E^{2}$ (2)
yields
$\frac{1}{E^{2}}=\big(\frac{2}{Tm^{2}c^{4}}\big)t\equiv Mt$ (3)
Eq. 3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the neutrinos will be found to lie on or near a straight line through the origin of slope M given by
$M=\frac{2}{Tm^{2}c^{4}}$ (4)
Thus, if the SN 1987A data show such a clustering about a straight line, from the best-fit slope M we can find the common mass of all neutrinos on or near the line as:
$m^{2}c^{4}=\frac{2}{TM}$ (5)
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line having negative slope.
To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin. Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions. It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the neutrino masses are too close to distinguish from one another.
Before we examine what the data show, there is a philosophical issue to mention. Given that in the standard view neutrino masses are less than $\Delta m^{2}_{atm}+\Delta m^{2}_{sol}$ – a value so small that any spread in neutrino travel times could not possibly be large enough to observe the neutrino mass, we have two choices on how to proceed. The first is not to even bother to see what the data suggest (the ”Cremonini option”) given that we know any positive result must be an anomaly, given the small upper limt on the neutrino mass . However, we have made the second choice, namely not to make any assumptions on the size of the neutrino masses and be guided by what the data themselves show.
IV Fitting 30 data points to 3 lines
Strikingly, when the SN 1987A data is plotted (see Fig.1) every one of the 30 neutrinos lies on one of three straight lines through the origin, and very few are next to two lines. The least square fit of the 30 data points to three lines through the origin has two free parameters. The first parameter is $S$ which adjusts the trial slopes of the three straight lines in Fig. 1 simultaneously and the second free parameter H is the size of the constant horizontal error bar on each data point in Fig.1. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square. However, only a few of the data points near the origin are within $1\sigma$ of more than one straight line. The good fit shown in Fig 1 has a chi square of 25.4. Given 28 degrees of freedom we find $p=61\%.$ The excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=18\%$.) For a fuller desccription of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig. 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot. Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is no more than 1-2 seconds.
Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to three straight lines that pass through the origin. Error bars in both horizontal and vertical are included in the fit. The best fit has $\chi^{2}=25.0$ for 28 dof, which yields $p=61\%$. The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq. 5. Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
The three masses corresponding to the slopes of the three straight lines are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.7eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line. The big difference from the author’s previously published values of this fit is in the value of $m_{2}$ because of the past omission of horizontal error bars, which are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line. We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.7eV/c^{2}$ and the $2.7eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN. (Serebrov2021, ; KATRIN2025, ). Of course, our $m_{1},m_{2},m_{3},$ however, are not sterile neutrino masses, but the three active neutrino masses.
Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points. Obviously for $H>>5$ sec error bars the fit has $p\approx 100\%$
V Discussion
In this section we discuss eleven issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply three specifc neutrino masses, and it is not just a mirage or background fluctuation.
V.1 Why not not exclude the LSD data?
Initially, the author rejected the LSD data. This rejection was not because it proviided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this and all the other reasons many people have for rejecting the LSD data can be answered, there really is no valid reason not to keep it.
V.2 Why are the LSD data bunched in time?
If the LSD burst of 5 neutrinos is genuine, it looks strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data has a kilosecond time scale. If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a bunch of events in a seconds-long time interval.
V.3 Why do the LSD neutrinos have the same E?
The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines. Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it. That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and further evidence related to the observed spectrum of MeV $\gamma$-rays from the galactic center. (Ehrlich2021, )
V.4 Alternate explanations for the LSD neutrinos?
All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
  • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}$
  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section 6.
  • If it was a genuine, low-energy pulse (e.g., from neutron capture on iron slabs around the detector), it would require the supernova to release an unusually high amount of energy in this form.
  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still creates the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
V.5 The meaning of the horizontal error bars?
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arriival time randomly shifted by a time up to that amount, and a horizonal error bar of that amount or greater would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 1-2 sec.
V.6 Does a $\Delta t=1-2$ sec agree with simulations?
We can find the actual size of the time window using a simulation of neutrino emission. In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 3. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background. These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 1-2 sec time window, and that our use of 2 sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
V.7 Why don’t late-emitted $\nu$’s mess up the fit?
First, because based on Fig.3 there are not very many of these, perhaps no more than 15-20% of the total for $t>2$ sec, and second because of the late neutrinos associated with the shallower sloped line in Fig. 1, a shift of the points horizontally even by quite a few seconds on a shallow sloped line has little effect on the goodness of its fit to the line, which depends mainly on the vertical error bar. Thus, for example, a data point lying one sigma away vertically from that line could still be less than one sigma away if the point was shifted as much as five seconds horizontally.
Figure 3: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the different time scales used for $t<1$ sec and $t>1$ sec. The fraction of the integrated luminosity for $t<1$ is less than it seems because of this scale change. Figure is used with permission of publisher
V.8 Two really big vertical error bars?
Two data points occurring at the largest t in Fig. 1 appear anomalous particularly because of their very large error bars compared to all others. This is the result of error propagation and the fact that on a plot of $1/E^{2}$ versus t the size of the vertical error bar varies as the inverse cube of the energy E according to $\Delta(1/E^{2})=2\Delta E/E^{3}.$ In any case these two points degrade the fit very little because of their large error bars.
V.9 What if all three $m_{k}$ are all very small?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.4. showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$ In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst. However, once again we would not be able to identify the three lines about which the 30 points are clustering, as long as we are assuming very small neutrino masses. This means that the double bang scenario can be ruled out, given Fig.1.
Figure 4: If the 3 neutrino masses were very small as in the standard neutrino model, and none had $m_{j}^{2}<0,$ the three lines on $1/E^{2}$ versus t plot about which the 30 data points would cluster might look like Fig. 4, at least for those neutrinos emitted exactly at two times one second apart. But since neutrinos are in fact emitted at a continuum of times from $t=0$ to $t=1$ sec, there are many such line triplets will be the lines about which data points would cluster. Given 1 or 2 sec horizontal error bars, it clearly would be impossible to find any clustering of points about 3 specific lines, as is actually the case in Fig.1.
V.10 The preposterous size of the three $m^{2}_{j}$
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data, whether it be comological data or data from the KATRIN experiment implying $m_{\nu}<1eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the effective mass has $m_{k}^{2}<0$ the effective mass can be arbitrarily close to zero or even negative. In order to look for the separate $m_{k}^{2}$ in a direct mass experiment like KATRIN one needs to combine three spectra one for each mass with appopriate weights (Ehrlich2019a, ). Likewise, the constraint imposed on the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ from cosmological data can be avoided if one or more of the tachyon masses is negative gravitationaly. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, so that they have a repulsive gravitational effect. The best upper limit on the total sum of neutrino masses from cosmology is $\sum m_{\nu}<0.12eV/c^{2}$ using data from the Planck satellite and the Atacama Cosmology Telescope. Without allowing negative $m_{k}$ all three would need to obey $m_{k}<0.12eV/c^{2},$ a limit that does not apply if negative $m_{k}$ are allowed. In light of these considerations, the numerical values from the three mass fit in Fig.1 are not necessarily ruled out by either lab or cosmological data.
V.11 The tail of the neutrino emission spectrum
It was noted earlier that background data not associated with the SN 1987A burst would not be expected to lie on three straight lines. Fortunately, Bozza et al. has obtained four Kamiokande II data points having times $15>t>25$ seconds on the day of SN 1989A. According to the “tail” of the emissions spectrum in Fig. 3, such data would have a negligible chance of being part of the SN 1987A burst, and they are likely to be background. Background events would be significantly lower in energy than supernova events. Because of that energy difference those four data points for $15>t>25$ seconds in fact have only a $1\%$ chance of being consistent with that best fit in Fig. 1 according to chi square.
VI what does it all mean?
The eleven considerations of the previous section suggest that the three neutrino masses found in the three line fit (Fig. 1) are real, even if one of the masses has $m^{2}<0.$ This is not the first time the author has claimed that the three neutrino masses can be found from SN 1987A data. When the evidence was first published in 2013, the three neutrino masses were interpreted in terms of a $3+3$ model, i.e., three active-sterile mass doublets, where each doublet was split by one of three oscillation $\Delta m^{2}.$ (Ehrlich2013, ; Ehrlich2019, ) We now discuss a much more plausible explanation involving CPT violation.
VI.1 Possible Violation of CPT Symmetry
CPT symmetry is at the heart of the standard model and is well-tested by looking for measured particle-antiparticle mass differences. CPT violation upper limits have been found for a number of pairs of particles and their antiparticles including proton-antiproton, electron-positron and $K_{0}-\bar{K}_{0}.$ Thus, for example $|m(K_{0})-m(\bar{K}_{0})|<0.50\times 10^{-18}GeV,$ which can also be expressed in dimensionless terms:
$X_{CPT}=\frac{|m(K_{0})-m(\bar{K}_{0})|}{m(K_{0})}<10^{-18}$ (6)
However, if viewed as a constraint on the mass-squared separation, the bound appears rather weak for kaons, namely only $0.25eV^{2}.$ It is generally believed that neutrinos might offer the best opportunity of any particle for observing CPT violation if it exists. The usual way of looking for violations or setting upper limits on CPT violation for neutrinos has not been to find a neutrino antineutrino mass difference, but rather to compare oscillation parameters for neutrinos and antineutrinos. There are currently two accelerator-based long baseline experiments T2K-II, NOvA-II and a reactor based medium-baseline JUNO experiment capable of testing CPT symmetry in this manner. The synergy of these three experiments should be able to set an upper bound in the $\Delta m^{2}$ between neutrinos and antineutrinos of $5.3\times 10^{-5}eV^{2}$, extending the current bound twenty fold. (Ngoc2023, ) Furthermore, an analysis of data from short-baseline experiments, combining Gallium radioactive source (neutrino) data and reactor (antineutrino) data, pointed to a possible CPT-violating asymmetry in the effective mixing angles with a statistical significance of about $3.5\sigma.$
The strongest hint of evidence for CPT violation has been provided by Hisakazu Minakata who notes that accelerator neutrino experiments in the $\nu_{\mu}\rightarrow\nu_{e}$ channel and the $\nu_{\mu}\rightarrow\nu_{\mu}$ channel show the problem of “appearance-disappearance tension” at a high confidence level. (Minakata2025, ) Minakata suggests non-unitarity as a simple and natural way of resolving the tension. He has constructed a model with a unique solution that is consistent with data, and implies a large neutrino-antineutrino asymmetry (hence CPT violation). Minakata’s model bridges the two highest CL signatures of CPT violation, i.e. data from the BEST and LSND-MiniBooNE experiments, and the model implies a large neutrino-antineutrino asymmetry.
As the preceding duscussion suggests, neutrinos are the ideal place to look for CPT violation because
  • Strong hints of CPT violation have been seen
  • The extent of the violation would be larger than for any other particles
  • The mechanism behind the tiny neutrino masses is not fully explained by the Standard Model.
  • Unlike charged particles, the weak neutrino interactions mean experiments are less susceptible to ”fake” CPT-violating signals that can arise from electromagnetic interactions.
  • Neutrinos can travel vast distances through matter without interacting significantly. Experiments use this to their advantage, studying neutrinos over very long baselines and at high energies.
A CPT violation might involve a mass difference between neutrino and antineutrino, which would bring about $\nu-\bar{\nu}$ oscillations. A relevant parameter specifying the magnitude of any CPT violation here is $(m^{2}_{\nu}-m^{2}_{\bar{\nu}})/m^{2}_{\nu}=\Delta m^{2}/m^{2}.$ Since there are three neutrino mass states one would expect that the CPT symmetry breaking parameter $X_{CPT}=\Delta m^{2}/m^{2}$ would be the same for the three masses, i.e.,
$X_{CPT}=\Delta m_{1}^{2}/m_{1}^{2}=\Delta m_{2}^{2}/m_{2}^{2}=\Delta m_{3}^{2}/m_{3}^{2}$ (7)
[…]
$X_{CPT}=\Delta m_{atm}^{2}/m_{1}^{2}=\Delta m_{sol}^{2}/m_{2}^{2}=\Delta m_{sbl}^{2}/m_{3}^{2}$ (8)
where we have used for $m_{1},m_{2}$ and $m_{3}$ in Eq. 7 the three masses found in our three mass fit to the SN 1987A data. Thus, we have the following dimensionless result for the amount of CPT violation from Eq. 8: $X_{CPT}=5\times 10^{-6},$ for the three mass states. Note that this postulated value is an order of magnitude smaller than that attainable in the three experiments T2K-II, NOvA-II and JUNO (which explains why no CPT violation has yet been directly seen).
The possible CPT violation represented by Eq. 8 may follow from a suggestion by Eichorn and Shiffer’s involving quantum gravity. (Eichorn2025, ) In the early universe during the transplankian regime when $\nu-\bar{\nu}$ oscillations did take place and the neutrino masses and their separations became frozen in. There is another implication of CPT violation for neutrinos involving tachyons. Various theoretical models of quantum gravity suggest it could lead to both CPT violation and modifications to the speed of neutrinos, resulting in subluminal and superluminal propagation, and so one might expect one or more member of the mass triplet defined by Eq. 5 be tachyonic. i.e., have $m^{2}_{j}<0.$ The recognition that neutrinos might provide evidence for CPT violation goes back as far as Bruno Pontecorvo (Pontecorvo1957, ) in 1957 who first suggested $\nu\bar{\nu}$ oscillations which can only occur if they have unequal masses.
VI.2 Why is CPT violation more likely than $3+3$?
The author previously postulated a $3+3$ model (3 active-sterile neutrino pairs) to explain the three outlandish masses found in the fit to the SN 1987A data (Ehrlich2019, ). There are many advantages of the CPT violation explanation over the earlier $3+3$ model. Thus, we have seen neutrinos are a natural place for CPT violation to occur, strong hints it is occurring, and there are difficulties with sterile neutrinos. The primary difficulty with sterile neutrinos is the contradiction with results from numerous other highly sensitive experiments which have found no evidence for them, as well as tension with cosmological data. In addition, with three sterile neutrinos and oscillations allowed between sterile and active neutrinos in the $3+3$ model one has a $6\times 6$ PMNS mixing matrix which appears to be in conflict with the standard well-established $3\times 3$ matrix parameters that are in good agreement with data.
Figure 5: An artists work depicting a 3D reconstruction based on SN 1987A observations. Credit ESO/L Calcada
VI.3 Summary and a conclusive test
This paper has made seven claims that some readers may consider very difficult to believe – see Table 1. Some of the seven claims will likely not be accepted as true without the strongest corroborating evidence.
1. No valid basis exists to ignore the LSD 5-hr early burst
2. No a priori upper limits exist on the three neutrino $m^{2}_{\nu,j}$
3. 3 $m_{j}^{2}$’s really can be found from 30 SN 1987A neutrinos
4. Negative search results cannot rule out tachyon’s existence
5. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. The three $m^{2}=2.7^{2}eV^{2},22.4^{2}eV^{2}$ and $2\times 10^{5}eV^{2}$
7. A likely explanation for these claims is CPT violation
Table 1: The seven claims made in this paper are based on the SN 1987A neutrino data. Regarding number 4, not only are negative search results incapable of proving tachyonic neutrinos non-existent, but they do not even necessarily make them less likely, but they may narrow the range of possible negative $m^{2}$ values.
We know SN 1987A was an asymmetric supernova (see Fig.5), with its expanding envelope and internal composition being lopsided rather than spherical. This spatial asymmetry was observed in the shape of the expanding envelope, and various other data. It is believed the asymmetry depicted likely arose from the progenitor star being a binary.
Figure 6: The asymmetric distribution of the debris of an explosion of the core of a 23-solar-mass progenitor star at 4.3 seconds after core bounce. The coloring codes mass density, blue being higher density. Credit: Argonne National Lab
Another particularly beautiful simulation of a supernova, this one lacking any symmetry whatsoever is shown in Fig. 6. However, it is not a mere spatial asymmetry that might corroborate the claims made in this paper, but rather a much deeper symmetry violation: the breaking of CPT symmetry and the evidence we have found supporting that violation by a dimensionless amount $X_{CPT}=5\times 10^{-5}.$ Usually, violations of the fundamental symmetries involve a decisive experiment. Thus, for parity (P) this was the C.S. Wu experiment, (Wu1957, ) for CP symmetry, this was the Fitch-Cronin experiment. (Fitch1964, ) Should a future experiment demonstrate a breaking of CPT symmetry for neutrinos by an amount $X_{CPT}=5\times 10^{-5}$ that would certainly be very strong corroboration of the seven claims. However, to find truly unambiguous evidence for all the claims, there is only one (and likely only one) decisive test, namely a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors.
The following concluding remarks represent one 88-year old’s supremely absurd optimism. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other six claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos, will definitely either confirm or reject the seven claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way to Earth from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say only 300 years). Were Betelgeuse to become a supernova, it would become incredibly bright, and even visible in daylight for many months. Go Betegeuse!
Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments.
The author would be incredibly grateful to any readers
of this paper who find errors or poorly explained sections.
Conflicts of interest
The author declares no conflicts of interest.