### v7

# Can the 3 neutrino masses really be found using SN 1987A data?

Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remarkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work. A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos. This absence of a sterile neutrino signal leads to two tests of the claim based on existing data in KATRIN and IceCube.
Keywords: SN 1987A, Neutrino, supernova, tachyon, KATRIN, IceCube, sterile neutrino
I Introduction
About $10^{58}$ neutrinos were emitted from the dying star Sanduleak 69 202 when it became the first supernova seen in 1987 (hence the name SN 1987A). Of those $10^{58}$ neutrinos only a few dozen were captured by Earth’s detectors. Yet, thousands of papers have been published about SN 1987A, and its well-studied nature is not surprising given that it is the only supernova in or close to our galaxy since neutrino detectors first existed. The earliest articles on the neutrino mass obtained from SN 1987A data, with one exception,(Cowsik1988, ) set only upper limits on its value. (Arnett1987, ; Burrows1988, ; Fang1987, ). The upper limits restriction is based on the almost universal belief that the values of the three neutrino mass states are simply too small to use the SN 1987A data to find them.
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. (KATRIN2025, ). This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation or rejection of the preposterous-sounding claim suggested in the title. The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
I.1 Are $m^{2}<0,$ and $v>c$ Tachyons “unphysical”?
Many theorists have regarded quantum fields giving rise to physical tachyons as being either impossible or unlikely in light of the many theoretical difficulties they create. However, unobservable $m^{2}<0$ particles seem to be OK with theorists. Thus, for most theorists the idea that symmetry-breaking in the interaction of Higgs fields with massless gauge fields leads to the production of unobservable tachyons is not objectionable. The difficulties with “physical” tachyons having an $m^{2}<0$ that can be observed include their energy spectrum being unbounded from below, their frame-dependent and unstable vacuum state, and their noncovariant commutation rules. Nevertheless, there are also theorists who are accepting of physical tachyons, and recently Paczos et al. (Paczos2024, ; Paczos2025, ) have shown that a covariant framework exists which allows for the proper quantization of tachyon fields which eliminates all such “unphysical” issues. Charles Schwartz is another theorist who has successfully challenged all the reasons why tachyons are considered unphysical, and shown how tachyons can account for many cosmological observations. (Schwartz2018, ; Schwartz2022, ).
I.2 Do negative searches make tachyons less likely?
Most reports of tachyons, including the 2011 OPERA experiment initial $v>c$ result, (Brumfiel2012, ) have later been shown to be incorrect apart from one-time sightings. (Ehrlich2022, ) Many physicists therefore believe that the tachyon hypothesis can be dismissed, or at least relegated to the highly implausible category. A more reasonable conclusion is that as long as not a single neutrino has ever been measured to have either $m^{2}>0$ or $v<c,$ the tachyonic neutrino hypothesis remains viable. A decline in belief in physical tachyons is perfectly understandable as a psychological reaction to negative results or even worse, false positive results like OPERA. However, negative and false positive results do not suggest tachyons do not exist, only that the $m^{2}$ for neutrinos is so close to zero, that we cannot yet tell if $m^{2}>0$ or $m^{2}<0$. You might think a negative hypothetical result $m^{2}=0.0\pm 0.1eV^{2}/c^{4}$ makes tachyonic neutrinos more unlikely than a previous result $m^{2}=0.0\pm 10.0eV^{2}/c^{4}$ because the allowed negative range in $m^{2}$ has been reduced a hundredfold. However, a Bayesian would realize that conclusion is false, because the allowed positive range has also been reduced by the same hundredfold amount. Neutrinos of course are the only known particles that might be tachyons since all others have been observed travelling at speeds $v<c,$ and tachyons if they exist must always be superluminal.
I.3 Assumptions in neutrino mass experiments
To date most experiments to measure the “effective mass” of the electron neutrino (or antineutrino) have done so by fitting the shape of the beta decay spectrum of tritium ${{}^{3}H}\rightarrow{{}^{3}He}+e+\bar{\nu_{e}}$ very close to its endpoint. It is assumed that the “effective mass” of the electron neutrino is defined in terms of the three states having masses $m_{j}$ and the PMNS mixing matrix $U_{i,j}$ as
$m_{\nu}^{2}=\Sigma|U_{i,j}|^{2}m_{j}^{2}$ (1)
For the electron, muon and tau neutrinos $i=1,2,3,$ respectively in Eq. 1. Many experiments in the past have reported negative values for $m_{\nu}^{2},$ for the electron neutrino, which were due to systematic effects, and not a true $m_{\nu}^{2}<0.$ The KATRIN experiment is the latest and most accurate experiment to measure $m_{\nu}^{2}$ and it has corrected for many of those earlier systematic effects, especially molecular final states. (Schneidewind2024, ) Thus, one assumption in direct mass experiments is that our current understanding of the molecular final states is accurately known. Would a future statistically significant $m^{2}<0$ mass then be attributed to a residual unknown systematic effect rather than an “unphysical” $m_{\nu}^{2}<0$ value? The use of the term unphysical for $m_{\nu}^{2}<0$ neutrinos in KATRIN publications suggests the former is more likely. (Drexlin2026, )
As of 2024 KATRIN has found only an upper limit to the mass, $m_{\nu}<0.45eV/c^{2}$ and by the end of the experiment they expect to reach a final sensitivity of better than $0.3eV/c^{2}$ at $90\%$ C.L. In addition, the best fit value as of 2024 is $m^{2}=-0.14^{+0.13}_{-0.15}eV^{2}.$ (KATRIN2024, ) This result, of course, leaves the true sign of $m^{2}$ ambiguous given the uncertainty. It is an interesting coincidence that this KATRIN value for $m_{\nu}^{2}$ is consistent with the predicted tachyonic mass the author made in a 2015 paper, i.e., $m^{2}=-0.11\pm 0.02eV^{2}.$ Ehrlich2015 While the statistical uncertainty on the $m^{2}$ value will be reduced as KATRIN takes more data, it might seem virtually certain that it will not shrink enough to eventually reveal its true sign to a $90\%$ C.L.
It is routinely claimed by KATRIN collaborators that direct mass experiments are “model-independent” and they make no assumptions (aside from energy conservation). (Drexlin2013, ; KATRIN2022, ; Parno2025, ) However, in addition to the several assumptions already noted, there is one other assumption routinely made. Namely, it is assumed that if the electron neutrino effective mass is close to zero, then the three masses $m_{j}$ contributing to it in Eq. 1 must also be close to zero, and hence not be experimentally resolvable. This assumption means that it is fine to fit the spectrum using a single mass. However, if one of the three masses satisfies $m_{j}^{2}<0,$ Eq. 1 would allow the three effective masses for $\nu_{e},\nu_{\mu}$ and $\nu_{\tau}$ to all be very close to zero, even when the individual $m_{j}$ are not. In fact, it will be shown using published neutrino data for SN 1987A that the three $m_{j}$ comprising the electron neutrino apparently have surprisingly large values.
I.4 Supernova SN 1987A neutrinos
Three neutrino detectors operated at the time supernova SN1987A occurred at 7:35 UT on Feb 23, 1987. Those detectors (Kamiokande II, IMB, and Baksan) recorded “bursts” or seconds-long groupings of 12, 8 and 5 antineutrinos respectively at approximately the same time, and they recorded the visible energy, incoming direction and arrival times of the neutrinos. The three detectors were not perfectly synchronized, with a time difference between them being perhaps a few seconds. This lack of synchronization is usually accounted for by assuming the first event in each detector occurred at time $t=0.$ The arriving electron antineutrinos were detected based on inverse beta decay $\bar{\nu_{e}}+p\rightarrow n+e^{+}$ so that the antineutrino energy was actually 1.3 MeV higher than the recorded visible $(e^{+})$ energy. Those three neutrino bursts occurred about three hours before the star emitting them suddenly brightened. This late light arrival, of course does not imply that neutrinos are superluminal. Rather, the three-hour gap between the neutrino burst and the later arrival of visible light, aligns with the small interaction between neutrinos with matter. Thus, photons but not neutrinos are delayed by interactions with the supernova ejecta.
I.5 The 5 hour early LSD neutrino burst
In addition to the three detectors, there was a fourth small detector known as LSD which was located under Mont Blanc. This detector was in fact better shielded from background radiation than the other three, and had a lower energy threshold, so in these two respects it was more sensitive. The LSD detector reported a burst of five neutrinos which occurred 484 min before the other three detectors’ received bursts. This early burst is therefore considered by most physicists to be a background fluctuation unrelated to SN 1987A because of its five hour early arrival. If the LSD burst were genuine it could constitute evidence that those neutrinos were tachyons, although other explanations also exist, including having SN1987A “bang” twice (Galeotti2016, ) and other ones as well. (Agafonova2018, ) Nevertheless, even raising the possibility of tachyons probably was enough to dismiss the existence of the LSD neutrino burst as unrelated to SN 1987A in the minds of most physicists. Moreover, there were four other reasons to be suspicious about the LSD burst besides the 5 hr early arrival, starting with it possibly being a statistical fluctuation of background.
The five neutrinos observed during a 7 second time interval in LSD would, if not connected to SN1987A, represent a very significant departure from background (roughly one neutrino every 80 seconds), and that background had been steady for months. In fact, the probability of a chance fluctuation giving rise to the LSD signal is extremely low, less than $1.4\times 10^{–6}.$ (Galeotti2016, ) Many observers also reject the LSD burst because no such early burst was seen in the three other detectors. However, the LSD neutrinos were all very low energy, and the other detectors could not have seen an LSD-like burst at that early time because of their higher energy thresholds. Thus, for example, the largest detector, Kamiokande II, had an energy threshold of 7 MeV, well above the energies of four of the five LSD neutrinos constituting the burst Another suspicious fact is that the main t=0 burst at 7:35 UT was not seen in LSD. However, that can also be explained because the small LSD detector did observe two neutrinos that time (not enough to qualify as a burst). Two final major problems with the LSD burst being associated with SN 1987A are (1) the very low energies of all the 5 neutrinos comprising it, and (2 ) their nearly equal energies or the virtual monochromaticity of the burst. In what follows we accept that the LSD neutrino burst was associated with SN 1987A, because all five objections to it are answerable, but we address the monochromaticity problem later.
II Finding $m^{2}$ of SN 1987A neutrinos
The possibility of using the SN 1987A data to find the three neutrino masses may seem implausible, but we shall see that it can be done if the neutrinos were emitted from the star close enough to simultaneously. Let us first see why for exactly simultaneous neutrino emissions the neutrino arrival time relative to a photon directly determines their travel time (relative to light), and how using the measured neutrino energy $E,$ and arrival time t we can compute the mass of the neutrinos. In what follows $T=168,000$ years is the time photons required to reach us from SN 1987A.
If a photon’s speed en route to Earth is $c=D/T$ then a neutrino’s speed would be $v=D/(t+T)\approx c(1-t/T).$ Here t is the time delay of a neutrino relative to a photon. But since there can also be such a delay if a neutrino is emitted later than other neutrinos, so the definition of t only makes sense if almost all neutrinos were emitted during a very small time window. For the neutrino speed we have v/c=1t/T.v/c=1-t/T. and also as an excellent approximation $v/c=\sqrt{1-m^{2}c^{4}/E^{2}}\approx 1-m^{2}c^{4}/2E^{2}$. Setting the two versions of $v/c$ equal
$v/c=1-t/T=1-m^{2}c^{4}/2E^{2}$ (2)
yields
$\frac{1}{E^{2}}=\big(\frac{2}{Tm^{2}c^{4}}\big)t\equiv Mt$ (3)
Eq. 3 implies that if a number of neutrinos with the same mass $m^{2}c^{4}$ were present and they had different arrival times t and energies E, then on a plot of $1/E^{2}$ versus $t,$ the data point for each neutrino will be found to lie on or near a straight line through the origin of slope M given by
$M=\frac{2}{Tm^{2}c^{4}}$ (4)
Thus, if the SN 1987A data show such a clustering about a straight line, from the best-fit slope M we can find the common mass of all neutrinos on or near the line as:
$m^{2}c^{4}=\frac{2}{TM}$ (5)
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin lying in the $t<0$ region.
To the extent that neutrinos all started out from SN 1987A at nearly the same instant, then since there are known to be three (and only three) non-sterile neutrino masses, every neutrino should lie on or close to one of three straight lines passing through the origin. Any neutrinos that do not do so must either be a background event unrelated to SN 1987A, or else the result of non-simultaneous emissions. It is however possible that the data points will all be found to be associated with only one or two straight lines rather than three lines if two or three of the lines (or neutrino masses) are too close to distinguish from one another.
III Fitting 37 data points to 3 lines
Thirty-seven neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events. Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event. (Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines. The least squares fit of 37 data points to three lines having a common $t=0$ origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig. 1 yielding 34 degrees of freedom. The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square.
III.1 The best fit
The best fit of 37 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin is shown in Fig. 1. Error bars in both horizontal and vertical are included in the fit. The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq. 5. Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity. The hard to see data point at t=2.5t=2.5sec carries a great deal of weight because of the small vertical error bar.The good fit shown in Fig 1 has a chi square of 29.9. Given 34 degrees of freedom we find $p=67\%.$ This good fit was for a horizontal error bar of $\pm 0.5$ seconds on all data points, and the fit would be much better for a $\pm 1.0$ second horizontal error bar ($p=91\%$.) For a fuller description of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig. 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot. Therefore the goodness of fits using 0.5 and 1 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is also around 0.5 to 1 seconds.
Figure 1: Best fit of 37 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin. See text and Table I for the meaning of the shaded region containing 7 “extra” Kamiokande II data points.
Seven data points in shaded region in Fig.1
time (sec) Hits     E(MeV)     $\Delta E$ (MeV)
13.7 30 10.8 2.1
15.1 31 11.2 2.1
15.4 25 9.0 1.9
16.5 26 9.4 1.9
18.9 32 11.5 2.1
18.9 28 10.1 2.0
22.8 33 11.9 2.2
Table 1: Seven “extra” Kamiokande II data points extracted from Fig. 4f of Hirata et al. (Hirata1988, ). These seven data points have $N_{hits}>24$ and $13<t<25$ seconds. The energy for each data point was obtained by the author based on the linear relation: $E=0.363N_{hits}$ to an accuracy of $5\%$ (Ehrlich2018, ). The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
III.2 The 7 “extra” Kamiokande II neutrinos
One problem in fitting the SN 1987A data to three straight lines is deciding what is the maximum time to use in the fit. This is important since after many seconds neutrinos are increasingly likely to be just background and not be associated with SN 1987A. Ref.  (Hirata1988, ) for Kamiokande II arbitrarily chose $t=12$ sec as the time cutoff. This choice resulting in their tabulating data only for 12 SN 1987A neutrinos. It was concerning to the author, however, that the last data point at $t=12$ seconds in Fig. 1 is $1.9\sigma$ above the fitted line. This concern motivated the desire to add more consecutive data points beyond $t=12$ sec, and seven of these were extracted by the author from ref. (Hirata1988, ) as explained in Table I. These data points are very unlikely to be due to background which lies below 20 hits, i.e. 20 photomultipliers activated, and adding them improved the goodness of the fit to three straight lines.
Based on Eq. 5 the three masses corresponding to the slopes of the three straight lines in Fig. 1 are $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ for the $m^{2}<0$ line. The author first discovered this three line fit to the SN 1987A data in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active-sterile neutrinos – see Fig.3.) There are two significant differences from the author’s previously published work on this fit (Ehrlich2013, ; Ehrlich2019, ). The first difference is in the use of data for Kamiokande II beyond the 12 neutrinos they tabulated, which slightly changes the value of $m_{1}.$ The second major difference is in the change in the value of $m_{2}$ because of the past omission of horizontal error bars. Those horizontal error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line. We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.10.
III.3 An omission in most sterile neutrino searches
The existence of sterile neutrinos is highly controversial, and there are probably more negative search results than positive ones. Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them. (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ). Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously, i.e., they do $3+1$ and not $3+2$ fits to data. There are two possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit, and another one being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
To date the only search for the pair of $m^{2}>0$ sterile neutrinos suggested by the SN 1987A data has been carried out by M. H. Chan and the author. (Chan2014, ) Sterile neutrinos are a natural possibility for the identity of dark matter in the galaxy and clusters of galaxies. Essentially, the authors first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies. Finally, the mass was varied to find the value that best fit the observed profile in each case. The results of these fits showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would indeed fit the galaxy clusters (lighter mass) and the Milky Way (heavier mass). This result offers a confirmation of the $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ masses deduced from the SN 1987A three mass fit.
Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 37 data points. Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>1.5$ sec error bars the fit probability $p\approx 100\%$
Figure 3: $3+3$ neutrino mass squared diagram postulated by the author in 2013, with updated $m^{2}$ values. There are 3 active-sterile doublets split by three $\Delta m^{2}$ for solar, atmospheric and short baseline (sbl) oscillations. Diagram is not to scale. The “haunted” area is for $m^{2}<0$ neutrinos. Many physicists regard $m^{2}<0$ tachyons as ghost particles. Illustration from Shutterstock.
IV Discussion
In this section we discuss 12 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0$ mass, really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
IV.1 Why not exclude the LSD data?
Initially, the author mistakenly rejected the LSD data. This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this issue and all the other reasons many people have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three active neutrinos the mandated number of straight lines about which clustering should occur is three. The burden of proof here should not be why not exclude this burst, but rather why not keep it – and there really is no compelling reason to ignore it.
IV.2 Why are the LSD data so bunched in time?
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data have a kilosecond time scale. If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst. The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they went backward in time, only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts. Because of that bunching at the left end of the LSD line the uncertainty in the slope (and the value of $m^{2}$) is $18\%$ or $m^{2}=-192,000\pm 34,000eV^{2}/c^{4}$.
IV.3 Why do the LSD neutrinos have the same E?
The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines. Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it. That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and support for the model from the observed spectrum of MeV $\gamma$-rays from the galactic center. (Ehrlich2021, )
IV.4 Alternate explanations for the LSD neutrinos?
Many phenomena in nature on the fringes of the known have alternative explanations. Most of the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus, perhaps the most likely alternate possibility is that a supernova can bang more than once. In fact, Galeotti and Pizzella have even claimed there was a third bang– not in the LSD detector but in Kamiokande II at a time about 20 min after the main burst. (Galeotti2007, ) Supernovae undergoing multiple bangs is a known phenomenon (known as Pulsational pair-instability supernovae), which can occur for very massive stars according to Woosley. (Woosley2017, ) However, the mass of SN 1987A was too small to experience the pair instability and in any case the phenomenon does not produce detectable neutrino bursts.
Other explanations of the LSD burst include:
  • If it was a statistical fluctuation it was a really big one $p=1.4\times 10^{-6}.$ In fact among the LSD signal candidates for the detection of a neutrino burst selected over 14 years of operation, it has the lowest background imitation probability.
  • If it was a genuine, low-energy pulse of neutrinos (e.g., from neutron capture on iron slabs around the detector), simulations show that their number is negligible. (Manukovskiy2022, ).
  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD burst.
  • The only other explanation known to this author for the LSD burst besides tachyons is the Froggatt-Nielsen proposal of “dark matter balls” inside the stellar core (Froggatt2015, ).
IV.5 Best size of the horizontal error bars?
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arrival time t randomly shifted to the right by a time up to that amount, and a horizonal error bar $H\approx\Delta t$ would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 0.5 to 1.0 seconds. Note that the time any individual neutrino is recorded is, of course, known to much greater precision than 0.5 seconds. However, the meaning of H takes into account the uncertainty of when the burst of neutrinos this neutrino was part of occurred, which is primarily determined by the choice of $t=0$ for the first arriving beutrino in a detector.
IV.6 A simulation of SN 1987A neutrino emissions
In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
These Bozza et al. results buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines. However, the use of as many as 9 free parameters in the Bozza spectral fit is a serious drawback, since with enough free parameters one can fit virtually any shape. In contrast, our fit to the $m^{2}>0$ neutrino spectrum used only two parameters, i.e., the two $m^{2}>0$ masses. Moreover, as long as the spread in neutrino emission times is small compared to the horizontal error bars used (as implied by Fig. 2), relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope, and there is no similar reqirement in the Bozza analysis.
Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the time scale change at $t=1$ seconds. The fraction of the integrated luminosity for $t<1$ seconds is only a bit more than half because of that scale change. Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
IV.7 Suppose neutrinos have $m<<1eV/c^{2}$?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines if emitted simultaneously, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we would not be able to find these lines because for 3 small masses they would lie on very many almost vertical lines filling the time window in Fig. 5. In effect each point lies on its own vertical line filling the time window. This is obviously far from what is the case for the actual data, which suggests that the neutrino mass does not satisfy $m<<1eV/c^{2}.$
Figure 5: How 50 neutrinos might appear on a $1/E^{2}$ versus t plot if their mass $m<<1eV/c^{2}$ and they were emitted during a time window of under 1 second. No congregation of the points about specific lines would be observed, completely unlike the real SN 1987A data which clearly clusters along three straight lines. Essentially, each of the 50 points lies on its own vertical line somewhere within the 1 second time window and no two points lie on the same line.
IV.8 The $\sum m_{\nu}$ constraint: a mass or flavor sum?
An apparent conflict exists between the three masses from the SN 1987A data fit and cosmological data. That conflict with the upper limit on the sum of the 3 neutrino masses $\Sigma m_{k}$ can be avoided, if one of the tachyon masses is negative gravitationally. (Schwartz2018, ) Thus, while tachyons are originally defined as having an imaginary kinematic mass, theorists often assume their gravitational mass is negative not imaginary, meaning they have a repulsive gravitational effect. The current upper limit on the total sum of neutrino masses from cosmology has tightened to $\sum m_{\nu}<0.07-0.12eV/c^{2}$ (Shao2025, ), but some analyses of the data even favor a negative value for the sum of the neutrino masses. (Elbers2025, ) There is an obvious conflict here between the current small limit on the sum and the three large masses found in the SN 1987A data, which could be avoided provided the mass sum were over flavor state masses not mass states.
In the recent literature the issue of whether the sum is over mass or flavor states is not widely discussed. However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states. Apart from historical usage there is a reason to view the sum as being over flavor state masses. Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old. At that time the two dominant processes capable of producing neutrinos were $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and both reactions would produce equal numbers of neutrinos for each flavor. For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each. Thus, the cosmological neutrino mass/energy density produced in that manner would be:
$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (6)
so, the appropriate constraint on R from fitting cosmological data is from Eq. 6 a constraint on the flavor state sum, and not the mass state sum, as sometimes assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case. Conversely, for the neutrino mass sum to be a sum over mass states, there would need to be a process that created neutrinos in equal numbers according to their mass state, and there is no such process.
IV.9 Consistency with direct mass experiments?
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass (see Eq.1) can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 best fit. In fact, given that the $3\times 3$ unitary mixing matrix has nine free parameters, all three effective masses could be close to zero.
IV.10 A “missing” $m_{4}=2.70eV/c^{2}$ sterile neutrino?
KATRIN recently reported finding no evidence for sterile neutrinos in its data based on the observed shape of the tritium beta decay spectrum. (KATRIN2025, ). It conducted this search by fitting its data to a sum of two spectra: that for a zero mass neutrino plus the spectrum for a sterile neutrino of mass $m_{4}$, which was assumed to take on various values and various weights $\alpha=sin^{2}2\theta.$ At each $(m_{4},\alpha)$ grid point, the $\chi^{2}$ function is minimized with respect to three free parameters. By examining the excluded regions of the plane based on the high $\chi^{2}$ KATRIN concluded there was no evidence for sterile neutrinos at a high confidence level in their data over the entire grid. This negative result was was in conflict with the Neutrino-4 group that had previously found evidence for a $2.70eV/c^{2}$ sterile neutrino in their oscillation experiment. Moreover, the conflict of KATRIN with the Neutrino-4 experiment was not a small one, since the data of the GALLEX, SAGE, and BEST experiments confirms the parameters of neutrino oscillations $\Delta m^{2}_{14}=7.3eV^{2}$ (or $m_{4}=2.70eV/c^{2}$ and $\sin^{2}2\theta_{14}=0.36)$ and the confidence level of the Neutrino-4 sterile neutrinos combined with the data of the GALLEX, SAGE, and BEST experiments is now $5.8\sigma.$ (Serebrov2024, )
IV.11 “The dog that didn’t bark” (the key section!)
It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, but here we consider a much more interesting possibility. Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos. However, KATRIN’s search was specifically looking for sterile neutrinos and it would not have worked for active ones as it was conducted. (KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig. 3), then KATRIN’s failure to see it is understandable because their search method was not appropriate for active neutrinos. Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. That’s because the zero mass effective mass is made up of the three specific masses we found for SN 1987A including one with $m^{2}<0.$ So, they should have fit their data using a combined differential spectrum $R_{C}$ for the three masses $m_{1}=2.70eV/c^{2},m_{2}=21.6eV/c^{2}$ and $m_{3}^{2}=-192,000eV^{2}/c^{4}$ with three adjustible weights, $\alpha_{j},$ that is:
$R_{C}=\sum\alpha_{j}R_{\beta}(E,E_{0},m^{2}_{\nu,j})+R_{Bkgd}$ (7)
where $R_{Bkgd}$ is the constant background rate, $R_{\beta}$ is the usual beta decay formula for a given $m^{2}$, E is the electron energy and $E_{0}$ is the fitted endpoint energy for tritium beta decay. For the weights $\alpha_{j}$ only two are independent since $\sum\alpha_{j}=1,$ and the $\alpha_{3}$ for the tachyonic mass is obviously very small in view of the very large magnitude of $|m_{3}^{2}|.$
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found. With the three masses set to their specified SN 1987A values (or else used as three pull terms in the fit), KATRIN should ideally observe a small concentrated area on a $(\alpha_{1},\alpha_{2})$ grid where good fits are found surrounded by a large area of poor fits. They could also do the same using other masses, where no good fits should be found. The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should also yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ KATRIN could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs. 93-94 in ref. (Rembielinski2021, ), as shown in their Fig. 1 the simple approach gives a result that is virtually indistinguishable from their full description.
In summary, if KATRIN were to redo its analysis with the preceding suggested changes it could be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$. The missing $2.70eV/c^{2}$ sterile neutrinos in the KATRIN search was the key clue here. It was effectively “the dog that didn’t bark,” a crucial example of a “negative fact.”
IV.12 Multiple sterile neutrinos in oscillation data?
The need for doing searches for multiple sterile neutrinos in a single search has already been noted in section III.3. How would one go about seeking evidence for both $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ and $m_{2}=21.6eV^{2}/c^{4}$ in oscillation data, bearing in mind that the oscillation wavelength for $m_{2}$ would be $21.6^{2}/2.7^{2}=64$ times shorter than the $m_{1}$ sterile neutrinos? It is unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were, one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see if there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},21.6^{2})eV^{2}/c^{2}.$ A more likely possibility would be the IceCube experiment. IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data. (Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$ To date, however, IceCube has only done a limited amount of searching for multiple sterile neutrinos. (Cabrera2024, )
V Summary and the ultimate test
This paper has made five claims that many readers may understandably consider very difficult to believe – see Table 2. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might convince many skeptics.
1. The LSD burst was due to SN 1987A but no 2nd bang
2. 3 $m_{j}^{2}$’s really can be found from 37 SN 1987A neutrinos
3. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
4. The three $m^{2}=2.7^{2}eV^{2},21.6eV^{2}$ and $1.92\times 10^{5}eV^{2}$
5. KATRIN and other data likely can confirm the above
Table 2: Five claims made in this paper.
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other four claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
These neutrinos will definitely either confirm or reject the five claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!
Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments. I also thank Guido Drexlin and Diana Parno of the KATRIN Collaboration for their kindness.
Conflicts of interest
The author declares no conflicts of interest.